OpenStudy (anonymous):

Answer This For a Medal ! Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5? (1, 7) (5, 0) (2, -1) (0, 5)

hero (hero):

y = y -3x + 5 = 2x + 5 5 - 5 = 2x + 3x 0 = 5x 0/5 = x 0 = x @gabbie96, If x = 0, can you find y ?